OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2)

OpenStudy (texaschic101):

find the slope first by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (4,1)..x1 = 4 and y1 = 1 (5,2)..x2 = 5 and y2 = 2 now we sub slope(m) = (2 - 1) / (5 - 4) slope(m) = 1 now we use y = mx + b form...we already found the slope(m), so now we will look for b, the y intercept. slope(m) = 1 you can use either of your sets of points...(4,1)..x = 4 and y = 1 now we sub y = mx + b 1 = 1(4) + b 1 = 4 + b 1 - 4 = b -3 = b our equation is : y = x - 3 but we need it in standard form : Ax + By = C y = x - 3 -- subtract x from both sides -x + y = -3 -- multiply by -1 to make x positive x - y = 3 <== standard form